GS stock is currently worth $56. Every year, it can increase by 30% or decrease by 10%. The stock pays no dividends, and the annual continuously-compounded risk-free interest rate is 4%. Using a two-period binomial option pricing model, find the price today of one two-year European put option on the stock with a strike price of $120.
Answer #1

The binomial option of option pricing use following formula in calculation as,

for PUT option , the pay off at last , Max [ (KSn), 0 ]

At pay off at previous node,

phpGXwkth.png

phpTMmhKB.png

U= Increase 30%= 1.30

d= decrease 10%= 0.90

P=\frac{e^{0.04*1}-0.90}{1.30-0.90}= 0.3521

U^2*S=1.312*56 94.64 25.36 max(120-94.64,0) U*S=1.3*56 72.8 42.49182 56 U*d*S=1.30*0.9*56 65.52 54.48 max(120-65.52,0) 54.768

Value at T=0 , $ 54.77